Question

Square Ltd’s projected sales report is given for the Month fromJuly to December 2017. Consider all sales would be credit sales. The actual sales for the months of May and June 2017 are also given.

Square Ltd Sales Report

For the year end 2017

May

June

July

August

September

October

November

December

$2,000

$5,890

$6,000

$13,200

$14,900

$22,000

$32,000

$22,600

Accounts Receivable is calculated at 50% current month, 30% last month and 20% second to last month.Build the projected cash collection for the last six months ending in December by using suitable mathematical tools.
Answer #1
Collection pattern:
Accounts receivable
50% of current month Hence,50% collected in the month of sale
30% of last month 50% collected in the first month of sale
Hence,collection in 2nd month=(100-50-30)%=20%
20% of second to last month 50% collected in the first month of sale
20% collected in the second month of sale
Hence,collection in 3rd month=(100-50-20-20)=10%'
Collection in last month Balance=(100-50-20-10)=20%
Collection pattern:
50% in the month of sale
20% in the second month of sales
10% in the third month of sales
20% in the fouth month of sales
Projected cash collections
July August September October November December Total
May sales 2000
10% in the third month of sales 200
20% in the fouth month of sales 400
June sales 5890
20% in the second month of sales 1178
10% in the third month of sales 589
20% in the fouth month of sales 1178
July sales 6000
50% in the month of sale 3000
20% in the second month of sales 1200
10% in the third month of sales 600
20% in the fouth month of sales 1200
August sales 13200
50% in the month of sale 6600
20% in the second month of sales 2640
10% in the third month of sales 1320
20% in the fouth month of sales 2640
September sales 14900
50% in the month of sale 7450
20% in the second month of sales 2980
10% in the third month of sales 1490
20% in the fouth month of sales 2980
October sales 22000
50% in the month of sale 11000
20% in the second month of sales 4400
10% in the third month of sales 2200
November sales 32000
50% in the month of sale 16000
20% in the second month of sales 6400
December sales 22600
50% in the month of sale 11300
Total 4378 8789 11868 16500 24530 22880 88945